• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 08:08:11 on October 14, 1998

    Entry number 2838

    keyword=Beam Time Accounting
    Hall A

    10/14/1998 0000-0800
    Experiment #: E94-010
    Shift Leader: slifer
    Shift Workers: T.Black, G. Kumbartzki, Z. Chai

    00000100020003000400050006000700Total Time
    Acceptable Beam
    in Use
    0 0 0 0 0 0 1 1 2
    Beam Available
    but Not in Use
    1 1 1 1 1 1 0 0 6
    Beam Not Available
    or Unacceptable
    0 0 0 0 0 0 0 0 0
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    0 0 0 0 0 0 1 1 2
    Planned Config-
    uration Change
    0 0 0 0 1 1 0 0 2
    Unplanned
    Experiment Down
    1 1 1 1 0 0 0 04
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 00:00- 04:00 repairs to target, E_Q3 quench reset. 04:00- 06:00 target NMR/EPR 06:00- 06:30 beam current optimization, Raster ,beam position 06:30- 08:00 Helium elastic (1.73 Gev) 9.4 M

    Data Entry by: slifer